OpenStudy (jmark):

If f(x) = logx^2(base2) and g(x) = log4(base x). Find the relation between f and g

OpenStudy (badhi):

\[g(x) = \log_x 4 = \frac{\ln 4}{\ln x}= 2\frac{\ln 2}{\ln x}\] \[f(x) = \log_2 x^2 = \frac{\ln x^2}{\ln 2} = 2\frac{\ln x}{\ln2}\] See a connection?